John is half as old as joke and joke is half as old as zainab the sum of their ages is 336years find their ages
Answer by colliefan(242)

If x = John's age:
Joke's age = 2x
Zainab's age = 2(2x)

The sum is 336 so:
x + 2x + 2(2x) = 336
x + 2x + 4x = 336
7x*1/7 = 336*1/7
x*1 = 336/7
x = 48
If x=48,
John is 48, Joke 96 and Zainab is 192.